Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability of marbles

Expert replies
by Stockmoose16 » Wed Sep 17, 2008 2:51 pm
If you have 4 red marbles, 3 yellow, and 2 green, what is the probability of getting 2 Red, 2 Green, and 1 yellow, if the marbles aren't replaced:

I tried this two ways and got two different answers:

METHOD #1: Combinations

# of ways to get 2R, 2G, 1Y 2: 4C2*2C2*3C1 =18
Total # of ways to pick 5 balls: 9C5= 126

Probability: 18/126

METHOD #2: Fractions

4/9*3/8*2/7*1/6*3/5 = 72/15120
Multiply by 5 for the # of possible rearrangements (or is it 5!... not sure)

= 360/15120

Why are these numbers not even close to each other?
Join the discussion
Source: — Problem Solving |

Re: Probability of marbles

by pseudononymous » Wed Sep 17, 2008 6:05 pm
Stockmoose16 wrote:If you have 4 red marbles, 3 yellow, and 2 green, what is the probability of getting 2 Red, 2 Green, and 1 yellow, if the marbles aren't replaced:

I tried this two ways and got two different answers:

METHOD #1: Combinations

# of ways to get 2R, 2G, 1Y 2: 4C2*2C2*3C1 =18
Total # of ways to pick 5 balls: 9C5= 126

Probability: 18/126

METHOD #2: Fractions

4/9*3/8*2/7*1/6*3/5 = 72/15120
Multiply by 5 for the # of possible rearrangements (or is it 5!... not sure)

= 360/15120

Why are these numbers not even close to each other?
I think method #1 is correct. The 2nd method restricts the balls to a specific order (red has to be picked first).
Join the discussion

Re: Probability of marbles

by lunarpower » Thu Sep 18, 2008 1:10 am
pseudononymous wrote:
Stockmoose16 wrote:If you have 4 red marbles, 3 yellow, and 2 green, what is the probability of getting 2 Red, 2 Green, and 1 yellow, if the marbles aren't replaced:

I tried this two ways and got two different answers:

METHOD #1: Combinations

# of ways to get 2R, 2G, 1Y 2: 4C2*2C2*3C1 =18
Total # of ways to pick 5 balls: 9C5= 126

Probability: 18/126

METHOD #2: Fractions

4/9*3/8*2/7*1/6*3/5 = 72/15120
Multiply by 5 for the # of possible rearrangements (or is it 5!... not sure)

= 360/15120

Why are these numbers not even close to each other?
I think method #1 is correct. The 2nd method restricts the balls to a specific order (red has to be picked first).
actually, both methods are correct. you just have to multiply by the correct number in method #2; that number is neither 5 nor 5!. instead, that number is the number of different ways of organizing two r's, two g's, and one y, which is 5! / (2!2!1!) , or 30.

i posted more about this problem here.
Ron has been teaching various standardized tests for 20 years.

--

Pueden hacerle preguntas a Ron en castellano
Potete chiedere domande a Ron in italiano
On peut poser des questions à Ron en français
Voit esittää kysymyksiä Ron:lle myös suomeksi

--

Quand on se sent bien dans un vêtement, tout peut arriver. Un bon vêtement, c'est un passeport pour le bonheur.

Yves Saint-Laurent

--

Learn more about ron
Join the discussion

by 4meonly » Thu Sep 18, 2008 6:17 am
somebode, what is OA?

I have 1/7

(4/9*3*8*2/7*1/6*3/5)*5!/2!*2! = 1/7
Join the discussion

by lunarpower » Fri Sep 19, 2008 2:28 am
4meonly wrote:somebode, what is OA?

I have 1/7

(4/9*3*8*2/7*1/6*3/5)*5!/2!*2! = 1/7
the answer to the problem is definitely 1/7, as mentioned in the thread linked above.
Ron has been teaching various standardized tests for 20 years.

--

Pueden hacerle preguntas a Ron en castellano
Potete chiedere domande a Ron in italiano
On peut poser des questions à Ron en français
Voit esittää kysymyksiä Ron:lle myös suomeksi

--

Quand on se sent bien dans un vêtement, tout peut arriver. Un bon vêtement, c'est un passeport pour le bonheur.

Yves Saint-Laurent

--

Learn more about ron
Join the discussion